### v1

@@ IV-C Discussion
Through two verifications, it was confirmed that the proposed method significantly improves the reproducibility of actions compared to MRS. MRS attempts to reproduce motion data based on the relationship between position and force determined by the original target’s impedance. In contrast, the proposed method recalculates position and force according to the current impedance, enhancing reproducibility. The rationale for this approach can be derived from the proposed method, as shown in (10)–(18). In the ideal case of position control, $x_{\mathrm{A}}$, $x_{\mathrm{B}}$, and $x^{\mathrm{cmd}}$ become equal. The same applies to velocity and acceleration. Consequently, according to (10)–(12) and (14), $x_{\mathrm{C}}$ becomes equal to $x^{\mathrm{cmd}}$. Moreover, by defining $den=f_{\mathrm{A}}-f_{\mathrm{B}}$ and substituting $x_{\mathrm{A}}=x_{\mathrm{B}}=x^{\mathrm{cmd}}$ into (16) and (17) and applying the result to (13) as:
@@
Equation (21) implies that if the motion data for Samples A and B were for position control, then the force command value $f_{\mathrm{C}}$ for Sample C is the force necessary to reach position $x^{\mathrm{cmd}}$, given the environmental impedance $m_{\mathrm{C}}$, $d_{\mathrm{C}}$, $k_{\mathrm{C}}$, $H_{\mathrm{C}}$. Similarly, for force control, $x_{\mathrm{C}}$ becomes equal to the position required for exerting $f_{\mathrm{C}}$ against the current environmental impedance. Therefore, it has been verified that the proposed method generates actions adapted to the current target by matching the impedance of the two motion data sets with the environmental impedance.
@@
Moreover, we confirmed the efficacy of the proposed method for samples with extrapolated impedance. Methods based on learning models usually interpolate the training data with approximations. Yet, beyond the scope of the training data, these methods might generate actions misaligned with user intentions in extrapolation scenarios. As described (21), the proposed method designs actions influenced by the current environmental impedance, thus facilitating adaptation to environments significantly divergent from the initial dataset. The uniform results demonstrate this adaptability for Patterns 2, 3, and 4, where extrapolation in both position and force control shows no notable variance from expected outcomes.
@@
This method generates actions based on two sets of preliminary data. In contrast, conventional machine learning-based motion generation methods require at least several dozen pieces of training data. This difference demonstrates that our method offers significant design, implementation, and practical advantages compared to machine learning approaches. However, machine learning methods include learning rough action plans (for example, recognizing the position of an object and moving forward or to the right accordingly) in the learning process. In contrast, our method does not undertake the generation of this planning part of the action. Therefore, it necessitates the design of a simple action planner. An MRS that compensates for variations in the location of objects based on image information has already been proposed[25, 26]. By integrating these, there is a potential to realize an MRS that adapts to objects’ location and physical properties. Our method focuses on simplifying the action correction part, and its usefulness in this field remains significant.
@@
Additionally, a significant advantage of the proposed method is its ability to generate actions without altering existing stable control systems. Traditional methods require adjustments to the control system, such as tuning the gains, making it essential to consider stability. However, our method assumes a stable control system, eliminating the need for additional discussions on stability. The only requirement for applying our method is the ability to measure and control position and force, making it easily applicable to existing manipulators and mobile units.
@@ IV-D Limitation
This paper estimated the impedance of the grasped object using multiple regression analysis with a 0.1 s time window, as described in Section II-C. In the current experimental environment, where the motor continuously performs grasping actions, sudden changes in impedance were not observed. However, in actual use environments, the repetition of contact and non-contact is expected, which means that changes in impedance could occur within the time window. If the impedance changes within the time window, it could lead to a decrease in the accuracy of the estimation. Therefore, exploring new methods for effectively estimating impedance within shorter time windows is necessary.
In this study, we demonstrated the efficacy of our proposed method through experiments using a single-degree-of-freedom rotary motor. In addition, we evaluated tasks in two distinct patterns: position control and force control. Given that this research is a foundational study, verification under simplified conditions was deemed necessary. The achievement in tasks exhibiting extreme impedance patterns is predicated on the assumption that similar outcomes can be anticipated for tasks with intermediate impedance values. Future work will focus on verification efforts using manipulators with multiple degrees of freedom and action data featuring intermediate impedance levels.
@@ V Conclusion
In this paper, we proposed a new method to enhance the adaptability of robot actions under diverse environmental conditions by matching the impedance of actions and the environment. This method is based on obtaining the impedance of actions from two sets of prior motion data and sequentially estimating the environmental impedance based on position and force information. Experiments conducted under conditions of differing action impedances, namely position control (with infinite impedance) and force control (with zero impedance), confirmed that the proposed method can reproduce actions adapted to the current environment.
The proposed method’s most significant advantage is its ability to generate effective actions with less data than machine learning-based methods. Additionally, it can be applied without changes to existing stable control systems, which is another significant benefit. These characteristics suggest that the proposed method has great potential in practical robot applications.
Future research will involve verification using motion data with intermediate impedance values. For example, this includes motion data where the impedance adjusts to grasp firmly when the object is rigid and gently when it is soft. Moreover, the focus will be on developing new estimation methods to effectively address changes in impedance in actions involving both contact and non-contact.

### v2

@@ IV-C Discussion
Through two verifications, it was confirmed that the proposed method significantly improves the reproducibility of actions compared to MRS. MRS attempts to reproduce motion data based on the relationship between the position and force determined by the original target’s impedance. In contrast, the proposed method recalculates the position and force according to the current impedance, thereby enhancing reproducibility. The rationale for this approach can be derived from the proposed method, as shown in (10)–(18). In the ideal case of position control, $x_{\mathrm{A}}$, $x_{\mathrm{B}}$, and $x^{\mathrm{cmd}}$ are equal. The same applies to velocity and acceleration. Consequently, according to (10)–(12) and (14), $x_{\mathrm{C}}$ is equal to $x^{\mathrm{cmd}}$. Moreover, by defining $den=f_{\mathrm{A}}-f_{\mathrm{B}}$ and substituting $x_{\mathrm{A}}=x_{\mathrm{B}}=x^{\mathrm{cmd}}$ into (16) and (17) and applying the result to (13) as:
@@
Equation (21) implies that if the motion data for Samples A and B are for position control, then the force command value $f_{\mathrm{C}}$ for Sample C is the force necessary to reach position $x^{\mathrm{cmd}}$, given the environmental impedances $m_{\mathrm{C}}$, $d_{\mathrm{C}}$, $k_{\mathrm{C}}$, and $H_{\mathrm{C}}$. Similarly, for force control, $x_{\mathrm{C}}$ is equal to the position required to exert $f_{\mathrm{C}}$ against the current environmental impedance. Therefore, it was verified that the proposed method generates actions adapted to the current target by matching the impedance of the two motion datasets with the environmental impedance.
@@
Moreover, we confirm the efficacy of the proposed method for samples with an extrapolated impedance. Methods based on learning models typically interpolate training data using approximations. However, beyond the scope of the training data, these methods may generate actions that are misaligned with user intentions in extrapolation scenarios. As described in (21), the proposed method designs actions influenced by the current environmental impedance, thus facilitating adaptation to environments that are significantly divergent from the initial dataset. The uniform results demonstrated this adaptability for Patterns 2, 3, and 4, where the extrapolation in both position and force control showed no notable variance from the expected outcomes.
@@
This method generates actions based on two sets of preliminary data. In contrast, conventional machine-learning-based motion generation methods require at least several dozen pieces of training data. This difference demonstrates that our method offers significant design, implementation, and practical advantages compared with machine-learning approaches. However, machine-learning methods include learning rough action plans (e.g., recognizing the position of an object and moving forward or to the right) during the learning process. In contrast, our method does not generate the planning part of the action. Therefore, it was necessary to design a simple action planner. An MRS that compensates for variations in the location of objects based on image information has been proposed [8, 31, 32]. By integrating these, an MRS that adapts to the location and physical properties of an object can be realized. Our method focuses on simplifying the action correction component; however, its usefulness in this field remains significant.
@@
Additionally, a significant advantage of the proposed method is its ability to generate actions without altering existing stable control systems. Traditional methods require adjustments to the control system, such as tuning the gains, making it essential to consider the stability. However, our method assumes a stable control system, thereby eliminating the need for additional stability discussions. The only requirement for applying our method is its ability to measure and control the position and force, making it easily applicable to existing manipulators and mobile units.
@@ IV-D Limitation
This study estimated the impedance of the grasped object using multiple regression analysis with a 0.1 s time window, as described in Section II-C. In the current experimental environment, where the motor continuously performed grasping actions, sudden changes in impedance were not observed. However, in actual use environments, the repetition of contact and noncontact is expected, which means that changes in impedance can occur within the time window. If the impedance changes within the time window, it could lead to a decrease in the estimation accuracy. Therefore, it is necessary to explore new methods for effectively estimating impedance within shorter time windows.
In this study, we demonstrated the efficacy of the proposed method through experiments using a rotary motor with a single degree of freedom (DOF). In addition, the tasks were evaluated using two distinct patterns: position control and force control. Given that this was a foundational study, verification under simplified conditions was necessary. The performance of tasks exhibiting extreme impedance patterns is predicted based on the assumption that similar outcomes can be anticipated for tasks with intermediate impedance values. Future work will focus on verification by using manipulators with multiple degrees of freedom and action data with intermediate impedance levels.
@@ V Conclusion
In this study, we propose a new method to enhance the adaptability of robot actions under diverse environmental conditions by matching the impedance of the actions and the environment. This method obtains the impedance of actions from two sets of prior motion data and sequentially estimates the environmental impedance based on position and force information. Experiments conducted under conditions of differing action impedances, namely, position control (with infinite impedance) and force control (with zero impedance), confirmed that the proposed method can reproduce actions adapted to the current environment.
The most significant advantage of the proposed method is its ability to generate effective actions with less data than machine-learning-based methods. In addition, it can be applied without changes to existing stable control systems. These characteristics suggest that the proposed method has great potential for practical robot applications.
Future research will involve verification by using motion data with intermediate impedance values. For example, this includes motion data in which the impedance is adjusted to grasp firmly when the object is rigid and gently when it is soft. Moreover, the focus will be on developing new estimation methods to effectively address changes in impedance in actions involving both contacts and non-contacts.